OpenStudy (anonymous):

@iambatman Can you please help me with this question? Given the exponential equation 3x = 243, what is the logarithmic form of the equation in base 10?

OpenStudy (freckles):

\[a^x=y \text{ is equivalent to } \log_{a}(y)=x\]

OpenStudy (anonymous):

Ok.so how do i plug it in?

OpenStudy (anonymous):

Help anyone???/

zepdrix (zepdrix):

\[\Large\rm 3^x=\color{orangered}{243}\]\[\Large\rm 3^x=\color{orangered}{3^?}\]243 is 3 to some power.